OpenStudy (anonymous):

will meadal Each marble bag sold by Soo's Marble Company contains 3 orange marbles for every 4 purple marbles. If a bag has 36 purple marbles, how many orange marbles does it contain?

OpenStudy (calculusxy):

we can set up a proportion here: \[\huge \frac{ orange }{ purple }\] we will always have to maintain that order in order to get the answer. what is also given is that the marbles are sold in a particular order: 3 orange marbles, for EVERY 4 purple marbles \[\large \frac{ 3orange}{ 4 purple }\] we know that there is 36 purple marbles, but we need to find the number of orange marbles. \[\large \frac{ 3orange }{ 4purple } = \frac{ x }{ 36 purple }\] as you can see, x represents the unknown number of orange marbles to solve for x, you need to "cross multiply" \[\large 36 \times 3 = 4x \]

OpenStudy (calculusxy):

solve for x

OpenStudy (anonymous):

27 ty

OpenStudy (calculusxy):

correct! yw :)